# Merging RLBWTs adaptively

### v5

@@ 1 Introduction
We describe here only how to merge the RLBWTs of two strings. Our algorithm can easily be extended to merge many RLBWTs or run-length compressed eBWTs, but we leave that to the full version of the paper. In Section 2 we review some preliminary concepts. As a warm-up, in Section 3 we discuss an easy algorithm for merging positional BWTs (PBWTs). In Section 4 we present our main algorithm and in Section 5 we present a first analysis that yields a time bound of $O((r\log r+L)\log(m+n))$. In Section 6 we give the proof of a technical lemma we use, whose details we refer to in Section 7 when we optimize our algorithm slightly to reduce the $r\log r$ in our time bound to $r$, so the whole bound becomes $O((r+L)\log(m+n))$.
@@ 2 Preliminaries
For the sake of brevity, we assume readers are familiar with run-length compression, the Burrows-Wheeler Transform (BWT), run-length compressed BWTs (RLBWTs), positional BWTs (PBWTs) [9], suffix arrays (SAs), the $\Psi$ function and longest common prefix (LCP) arrays; see Mäkinen et al.’s [17] and Navarro’s [19] texts for an introduction. The extended BWT (eBWT) [18] of two strings $S[1..m]$ and $T[1..n]$ is an interleaving $\mathrm{BWT}_{S,T}[1..m+n]$ of the characters in the BWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..m]$ of $S$ and $T$. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ are subsequences of $\mathrm{BWT}_{S,T}$, if it has $r$ runs then they each have at most $r$ runs.
@@ 3 Merging PBWTs
Because the $j$th columns of the two input PBWTs already contain all those next bits and in the correct order, we can use the input PBWTs instead of the explicit haplotypes. To do this, we start with an empty draft of the $j$th column of the merged PBWT and scan the (j1)(j-1)st column of the merged PBWT twice. During the first scan, whenever we see a 0 we copy a bit (the next bit in that 0’s haplotype) from the $j$th column of the input PBWT that is the source of that 0. During the second scan, whenever we see a 1 we copy a bit (the next bit in that 1’s haplotype) from the $j$th column of the input PBWT that is the source of that 1. However, we may be able to work more efficiently considering blocks rather than literally bit by bit.
@@
Theorem 3.1 may be of independent interest but we have proven it mainly as a warm-up for merging RLBWTs, which is significantly more complicated.
@@ 4 Main algorithm
Suppose we are to merge the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of strings $S[1..m]$ and $T[1..n]$ into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ of $S$ and $T$ together, and we have already built the data structures for $\Psi$ for $S$ and $T$ described in Lemma 2.1 so we can compare characters’ contexts. The pseudocode for our algorithm is shown in Figure 2.
@@
By the definition of the eBWT, it is impossible for $\mathrm{BWT}_{S}[i+k]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and for $\mathrm{BWT}_{T}[j+\ell]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$ in $\mathrm{BWT}_{S,T}$, since then $\mathrm{BWT}_{S}[i+k]$ would have to both precede and follow $\mathrm{BWT}_{T}[j+\ell]$. Therefore, the only remaining case with $\mathrm{BWT}_{S}[i]=\mathrm{BWT}_{T}[j]$ is when either $i+k=m+1$ or $\mathrm{BWT}_{S}[i+k]$ follows the entire run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and either $j+\ell=n+1$ or $\mathrm{BWT}_{T}[j]$ follows the entire run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$. In this case, we output
@@ 5 First analysis
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r\log r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of its irreducible LCP values.
@@ 7 Optimization
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of its irreducible LCP values.
@@
Let $P_{j}$ and $Q_{j}$ be the contents of $T_{P}$ and $T_{Q}$ after $j$ rounds, so $P_{0}=P$ and $Q_{0}=Q$. Suppose we remove $(s,e)$ from $L$ during the (j+1)(j+1)st round, meaning $s,e)\cap P_{j}\geq 2d$. We use $T_{P}$ to find the (d+1)(d+1)st largest element $p$ in $s,e)\cap P_{j}$. We use $T_{Q}$ to find $\pi^{-1}(p)$ by finding the predecessor $q$ of $p$ in $Q_{j}$ and adding pqp-q to $\pi^{-1}(q)$, which is stored as satellite data with $q$. We insert $\pi^{-1}(p)$ into $T_{P}$ with $p$ as satellite data, and insert $p$ into $T_{Q}$ with $\pi^{-1}(p)$ as satellite data. Therefore $Q_{j+1}=\{\pi(i)\ :\ i\in P_{j+1}\}\cup\{n+1\}$.

### v6

@@ 1 Introduction
We describe here only how to merge the RLBWTs of two strings. Our algorithm can easily be extended to merge many RLBWTs or run-length compressed eBWTs, but we leave that to the full version of the paper. In Section 2 we review some preliminary concepts. As warm-ups, in Section 3 we present simple algorithms for merging positional BWTs (PBWTs) and RLBWTs. In Section 4 we present our main algorithm and in Section 5 we present a first analysis that yields a time bound of $O((r\log r+L)\log(m+n))$. In Section 6 we give the proof of a technical lemma we use, whose details we refer to in Section 7 when we optimize our algorithm slightly to reduce the $r\log r$ in our time bound to $r$, so the whole bound becomes $O((r+L)\log(m+n))$.
@@ 2 Preliminaries
For the sake of brevity, we assume readers are familiar with run-length compression, the Burrows-Wheeler Transform (BWT), run-length compressed BWTs (RLBWTs), positional BWTs (PBWTs) [9], suffix arrays (SAs), the $\Psi$ function and longest common prefix (LCP) arrays; see Mäkinen et al.’s [17] and Navarro’s [19] texts for an introduction. The extended BWT (eBWT) [18] of two strings $S[1..m]$ and $T[1..n]$ is an interleaving $\mathrm{BWT}_{S,T}[1..m+n]$ of the characters in the BWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of $S$ and $T$. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ are subsequences of $\mathrm{BWT}_{S,T}$, if it has $r$ runs then they each have at most $r$ runs. For convenience, we write $\mathrm{BWT}_{S}$, $\mathrm{BWT}_{T}$ and $\mathrm{BWT}_{S,T}$ to denote both the BWTs and the RLBWTs, stating when they are run-length compressed.
@@ 3 Warm-ups
Because the $j$th columns of the two input PBWTs already contain all those next bits and in the correct order, we can use the input PBWTs instead of the explicit haplotypes. To do this, we start with an empty draft of the $j$th column of the merged PBWT and scan the (j1)(j-1)st column of the merged PBWT twice. During the first scan, whenever we see a 0 we copy a bit (the next bit in that 0’s haplotype) from the $j$th column of the input PBWT that is the source of that 0. During the second scan, whenever we see a 1 we copy a bit (the next bit in that 1’s haplotype) from the $j$th column of the input PBWT that is the source of that 1. However, our algorithm may work more efficiently considering blocks rather than literally bit by bit.
@@
[…]
Suppose we are to merge the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of strings $S[1..m]$ and $T[1..n]$ into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ of $S$ and $T$ together, and we have already built the data structures for $\Psi$ for $S$ and $T$ described in Lemma 2.1 so that we can compare characters’ contexts. For example, if $S$ is the concatenation of the \$-terminated strings GATTACAT, GATACAT and GATTAGATA and $T$ is the concatenation of those three strings’ \$-terminated reverse complements, then
$\mathrm{BWT}_{S}$ $=$ $\mathrm{TAT^{5}C^{2}G^{4}A^{3}\mathdollar^{3}A^{3}TATA^{2}}$
$\mathrm{BWT}_{T}$ $=$ $\mathrm{C^{3}T^{2}A^{2}T^{2}\mathdollar^{2}T^{6}CG^{2}\mathdollar A^{6}}$
$\mathrm{BWT}_{S,T}$ $=$ $\mathrm{C^{2}TATCT^{6}C^{2}G^{2}A^{2}T^{2}\mathdollar^{2}G^{2}T^{3}A^{2}TA\mathdollar^{3}T^{2}A^{3}CGTATG\mathdollar A^{8}}\,$,
with exponents indicating run lengths, and $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ are interleaved in $\mathrm{BWT}_{S,T}$ as shown in Figure 2. In this toy example, $r=27$, $b=18$, $B=18$, $\sigma=5$ and $m+n=54$.
0 \$ATGTAATC\$ATGTATC\$TATCTAAT C
\$ATGTATC\$TATCTAATC\$ATGTAAT C
1 \$GATACAT\$GATTAGATA\$GATTACA T
\$GATTACAT\$GATACAT\$GATTAGAT A
\$GATTAGATA\$GATTACAT\$GATACA T
1 \$TATCTAATC\$ATGTAATC\$ATGTAT C
0 A\$GATTACAT\$GATACAT\$GATTAGA T
1 AATC\$ATGTAATC\$ATGTATC\$TATC T
AATC\$ATGTATC\$TATCTAATC\$ATG T
1 ACAT\$GATACAT\$GATTAGATA\$GAT T
ACAT\$GATTAGATA\$GATTACAT\$GA T
AGATA\$GATTACAT\$GATACAT\$GAT T
AT\$GATACAT\$GATTAGATA\$GATTA C
AT\$GATTAGATA\$GATTACAT\$GATA C
ATA\$GATTACAT\$GATACAT\$GATTA G
ATACAT\$GATTAGATA\$GATTACAT\$ G
2 ATC\$ATGTAATC\$ATGTATC\$TATCT A
ATC\$ATGTATC\$TATCTAATC\$ATGT A
ATC\$TATCTAATC\$ATGTAATC\$ATG T
ATCTAATC\$ATGTAATC\$ATGTATC\$ T
ATGTAATC\$ATGTATC\$TATCTAATC \$
ATGTATC\$TATCTAATC\$ATGTAATC \$
2 ATTACAT\$GATACAT\$GATTAGATA\$ G
ATTAGATA\$GATTACAT\$GATACAT\$ G
0 C\$ATGTAATC\$ATGTATC\$TATCTAA T
C\$ATGTATC\$TATCTAATC\$ATGTAA T
C\$TATCTAATC\$ATGTAATC\$ATGTA T
1 CAT\$GATACAT\$GATTAGATA\$GATT A
CAT\$GATTAGATA\$GATTACAT\$GAT A
1 CTAATC\$ATGTAATC\$ATGTATC\$TA T
0 GATA\$GATTACAT\$GATACAT\$GATT A
GATACAT\$GATTAGATA\$GATTACAT \$
GATTACAT\$GATACAT\$GATTAGATA \$
GATTAGATA\$GATTACAT\$GATACAT \$
1 GTAATC\$ATGTATC\$TATCTAATC\$A T
GTATC\$TATCTAATC\$ATGTAATC\$A T
0 T\$GATACAT\$GATTAGATA\$GATTAC A
T\$GATTAGATA\$GATTACAT\$GATAC A
TA\$GATTACAT\$GATACAT\$GATTAG A
2 TAATC\$ATGTAATC\$ATGTATC\$TAT C
TAATC\$ATGTATC\$TATCTAATC\$AT G
2 TACAT\$GATACAT\$GATTAGATA\$GA T
TACAT\$GATTAGATA\$GATTACAT\$G A
TAGATA\$GATTACAT\$GATACAT\$GA T
2 TATC\$TATCTAATC\$ATGTAATC\$AT G
TATCTAATC\$ATGTAATC\$ATGTATC \$
TC\$ATGTAATC\$ATGTATC\$TATCTA A
TC\$ATGTATC\$TATCTAATC\$ATGTA A
TC\$TATCTAATC\$ATGTAATC\$ATGT A
TCTAATC\$ATGTAATC\$ATGTATC\$T A
TGTAATC\$ATGTATC\$TATCTAATC\$ A
TGTATC\$TATCTAATC\$ATGTAATC\$ A
1 TTACAT\$GATACAT\$GATTAGATA\$G A
TTAGATA\$GATTACAT\$GATACAT\$G A
Figure 2: The lexicographically sorted cyclic shifts of the concatenation of the three \$-terminated strings GATTACAT, GATACAT and GATTAGATA (in red), and of the concatenation of those three strings’ \$-terminated reverse complements (in blue), with the LCP values shown at block boundaries (in black). The eBWT of the two concatenations is the last column of the matrix, slightly set from the rest. The red subsequence of the eBWT is the BWT of the first concatenation and the blue subsequence is the BWT of the second.
We first set $i$ and $j$ to 1, then compare $\mathrm{context}_{S}(i)$ and $\mathrm{context}_{T}(j)$ to check whether the first block of characters in $\mathrm{BWT}_{S,T}$ is from $\mathrm{BWT}_{S}$ or $\mathrm{BWT}_{T}$, and set a Boolean flag to true in the former case and false in the latter. As long as $i\leq m$ and $j\leq n$, if the flag is true then we use a doubling search in $\mathrm{BWT}_{S}[i..m]$ to find the largest value $i^{\prime}$ with $\mathrm{context}_{S}(i^{\prime})\prec\mathrm{context}_{T}(j)$, output $\mathrm{BWT}_{S}[i..i^{\prime}]$, reset $i$ to $i^{\prime}+1$, and reset the flag to false. If the flag is false then we use a doubling search in $\mathrm{BWT}_{T}[j..n]$ to find the largest value $j^{\prime}$ with $\mathrm{context}_{T}(j^{\prime})\prec\mathrm{context}_{S}(i)$, output $\mathrm{BWT}_{T}[j..j^{\prime}]$, reset $j$ to $j^{\prime}+1$, and reset the flag to true. Eventually, when $i=m+1$ or $j=n+1$, we output either $\mathrm{BWT}_{T}[j..n]$ or $\mathrm{BWT}_{S}[i..m]$, respectively. The pseudocode for our algorithm is shown in Figure 3.
$1i\leftarrow$ 1
[…]
3:if $\mathrm{context}_{S}(i)\prec\mathrm{context}_{T}(j)$ then
4:  $\mathit{flag}\leftarrow\mathrm{true}$
5:else
6:  $\mathit{flag}\leftarrow\mathrm{false}$
7:end if
8:while $i\leq m$ and $j\leq n$ do
9:  if $\mathit{flag}$ then
10:   doubling search for the largest value $i^{\prime}$ with $\mathrm{context}_{S}(i^{\prime})\prec\mathrm{context}_{T}(j)$
11:   output $\mathrm{BWT}_{S}[i..i^{\prime}]$
12:   $i\leftarrow i^{\prime}+1$
13:   $\mathit{flag}\leftarrow\mathrm{false}$
14:  else
15:   doubling search for the largest value $j^{\prime}$ with $\mathrm{context}_{T}(j^{\prime})\prec\mathrm{context}_{S}(i)$
16:   output $\mathrm{BWT}_{T}[j..j^{\prime}]$
17:   $j\leftarrow j^{\prime}+1$
18:   $\mathit{flag}\leftarrow\mathrm{true}$
19:  end if
20:end while
21:if $\mathit{flag}$ then
22:  output $\mathrm{BWT}_{S}[i..m]$
23:else
24:  output $\mathrm{BWT}_{T}[j..n]$
25:end if
By induction, during each pass through the while loop, we output the next block of $\mathrm{BWT}_{S,T}$, from $\mathrm{BWT}_{S}$ if the flag is true and from $\mathrm{BWT}_{T}$ if the flag is false. If the flag is true then all of the $O(\log m)$ comparisons we perform during the current pass are against $\mathrm{context}_{T}(j)$, which is in the block from $\mathrm{BWT}_{T}$ we will output during the next pass, so the number of characters we extract for each of those comparisons is at most 1 plus the maximum of the LCP values at the boundaries of that block from $\mathrm{BWT}_{T}$. If the flag is false then all of the $O(\log n)$ comparisons we perform during the current pass are against $\mathrm{context}_{S}(i)$, which is in the block from $\mathrm{BWT}_{S}$ we will output during the next pass, so the number of characters we extract for each of those comparisons is at most 1 plus the maximum of the LCP values at the boundaries of the block from $\mathrm{BWT}_{S}$.
[…]
Theorem 3.2.
[…]
@@ 4 Main algorithm
Again, suppose we are to merge the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of strings $S[1..m]$ and $T[1..n]$ into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ of $S$ and $T$ together, and we have already built the data structures for $\Psi$ for $S$ and $T$ described in Lemma 2.1 so that we can compare characters’ contexts. The pseudocode for our algorithm is shown in Figure 4.
@@
By the definition of the eBWT, it is impossible for $\mathrm{BWT}_{S}[i+k]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and for $\mathrm{BWT}_{T}[j+\ell]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$ in $\mathrm{BWT}_{S,T}$, since then $\mathrm{BWT}_{S}[i+k]$ would have to both precede and follow $\mathrm{BWT}_{T}[j+\ell]$. Therefore, the only remaining case with $\mathrm{BWT}_{S}[i]=\mathrm{BWT}_{T}[j]$ is when either $i+k=m+1$ or $\mathrm{BWT}_{S}[i+k]$ follows the entire run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and either $j+\ell=n+1$ or $\mathrm{BWT}_{T}[j+\ell]$ follows the entire run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$. In this case, we output
@@ 5 First analysis
Given the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r\log r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of its irreducible LCP values.
@@ 7 Optimization
Given the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of its irreducible LCP values.
With care, it is possible to combine Theorems 3.2 and 7.1 and obtain the advantages of each — better than simply dovetailing — but, due to space constraints, we leave that as an exercise for the reader until we publish the full version of this paper.
@@
Let $P_{j}$ and $Q_{j}$ be the contents of $T_{P}$ and $T_{Q}$ after $j$ rounds, so $P_{0}=P$ and $Q_{0}=Q$. Suppose we remove $(s,e)$ from $L$ during the (j+1)(j+1)st round, meaning $s,e)\cap P_{j}\geq 2d$. We use $T_{P}$ to find the (d+1)(d+1)st smallest element $p$ in $s,e)\cap P_{j}$. We use $T_{Q}$ to find $\pi^{-1}(p)$ by finding the predecessor $q$ of $p$ in $Q_{j}$ and adding pqp-q to $\pi^{-1}(q)$, which is stored as satellite data with $q$. We insert $\pi^{-1}(p)$ into $T_{P}$ with $p$ as satellite data, and insert $p$ into $T_{Q}$ with $\pi^{-1}(p)$ as satellite data. Therefore $Q_{j+1}=\{\pi(i)\ :\ i\in P_{j+1}\}\cup\{n+1\}$.